OpenStudy (anonymous):

Find the volume of the region lying below z = 8 - x2 - y2 and above the xy-plane.

OpenStudy (anonymous):

it is calc 3! double integral i think!

OpenStudy (kinggeorge):

First, you have to find where \(z=8-x^2-y^2\) intersects the plane \(z=0\). From there you can find the bounds with which to integrate.

OpenStudy (anonymous):

i know that is the general scheme, but i cant set it up so if you can thatd be sweet bro

OpenStudy (kinggeorge):

Well, if \(z=0\), then \(x^2+y^2=8\). In other words, this is a circle of radius \(\sqrt{8}\) centered at the origin.

OpenStudy (kinggeorge):

Thus, we might want to switch to polar coordinates. Have you worked with polar yet?

OpenStudy (anonymous):

yes i was just about to say that, yes i have worked with them how would that look?

OpenStudy (kinggeorge):

You would want to integrate r first, and then \(\theta\). Also, you know that \(x^2+y^2=r^2\) and \(\theta\) is the full rotation around. So the integral should look something like\[\int\limits_0^{2\pi} \int\limits_0^{\sqrt{8}} r(8-r^2) \quad dr \;\;d\theta\]

OpenStudy (anonymous):

wow man great job! thanks alot!

OpenStudy (kinggeorge):

You're welcome.